OpenStudy (anonymous):

which list shows all the positive factors of 15? 1,15 1,5,15 1,3,15 1,3,5,15 last question!!

OpenStudy (chriss):

1,3,5,15

OpenStudy (anonymous):

study up your multiplication tables and these will come easier to you.

OpenStudy (anonymous):

i know but its really hard to rememeber:\

OpenStudy (anonymous):

it's a pain at first but you will find everything else for many subjects so much easier once you get them down.

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

Here are some tricks to better remember your multiplication tables. 2: Are always even. 2*8 = 16 5: Always end in 5 or 0. 5*5 = 25 & 5*2 = 10 6: Always also divisible by 3 9: Digits always add to 9 when 2 digits. 4*9 = 36; 3+6 = 9 10: Easy enough. Just add a zero. 11: Digits repeat when 2 digits. 2*11 = 22; Middle digit is one plus last digit when 3 digits. 11*11 = 121; 11*12 = 132 Refer here for further explanation: http://mathforum.org/k12/mathtips/multiplication.tips.html